# Computing the $q$-Multiplicity of the Positive Roots of $\mathfrak{sl}_{r+1}(\mathbb{C})$ and Products of Fibonacci Numbers

### v1

@@ 2 Weyl Alternation Sets
We proceed by induction. For $n=1$, the only options are to select no integers, hence the $\emptyset$, or to select $1$. There are two such choices and note $F_{1+2}=F_{3}=2$, and more over $F_{3}=F_{2}+F_{1}$, as claimed. Assume that for all $k\leq n$ the number of ways to select a subset of $[k]$ consisting of only nonconsecutive integers is given by $F_{k+2}$. Now let $k=n+1$. Note that for any subset of $[n-1]$ consisting of nonconsecutive integers, we can include the integer $n+1$ to get a subset of $[n+1]$ consisting of nonconsecutive integers. Then by induction, there are $F_{n-1+2}=F_{n+1}$ many such subsets, all of which contain the integer $n+1$. Moreover, any subset of $[n]$ consisting of nonconsecutive integers is also a subset of $[n+1]$ with the same property and do not contain $n+1$. By induction hypothesis, there are $F_{n+2}$ such subsets. As these subsets are disjoint (since they either contain $n+1$ or not), then we have shown that the number of subsets of $[n+1]$ consisting of nonconsecutive integers is given by $F_{n+2}+F_{n+1}$, which is equal to $F_{n+3}$, as desired. ∎
Harris in [7] gives another enumerative result which we use in the following section counts subsets of $k$ nonconsecutive integers from the set $[n]$.
@@
If $1\leq i\leq j\leq r$, then
@@
Selecting a subset consisting of nonconsecutive integers from the set $\{2,\ldots,i-1\}$ is the same as selecting a subset consisting of nonconsecutve integers from the set $\{1,\ldots,i-2\}$. By Proposition 2.3, there are $F_{i}$ such subsets. The same process can be done for selecting a subset consisting of nonconsecutve integers from the set $\{j+1,\ldots,r-1\}$ to get $F_{r-j+1}$ such subsets.
The choices for subsets $\{2,\ldots,i-1\}$ and $\{j+1,\ldots,r-1\}$ are disjoint and independent of each other. Thus the total number of ways to select nonconsecutive integers is given by the product $F_{i}\cdot F_{r-j+1}$. Therefore, $|\mathcal{A}(\tilde{\alpha},\alpha_{i}+\cdots+\alpha_{j})|=F_{i}\cdot F_{r-j+1}$. ∎
@@ 3 The $q$-Analog of Kostant’s Partition Function
[…]
@@
[…]
@@
Let $\beta=\alpha_{1}+\alpha_{2}+\cdots+\alpha_{s}$ and $\gamma=\alpha_{t}+\alpha_{t+1}+\cdots+\alpha_{s+t-1}$. Notice that $h(\beta)=s$ and $h(\gamma)=(s+t-1)-t+1=s$. By [6, Proposition 3.1], we know
@@
The following propositions give the value of the $q$-analog of Kostant’s partition function when the input is $\sigma(\tilde{\alpha}+\rho)-\rho-\mu$, with $\mu\in\Phi^{+}$ and $\sigma\in\mathcal{A}(\tilde{\alpha},\mu)$.
@@
    $\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$ =$=\begin{cases}q^{\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)}&\mbox{if $s_{j+1}$ is contained in $\sigma$}\\ q^{\ell(\sigma)+1}(1+q)^{r-h(\mu)-2\ell(\sigma)-1}&\mbox{otherwise.}\end{cases}$ (9)
@@
    $\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$ =$=\begin{cases}q^{\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)}&\mbox{if $s_{i-1}$ is containted in $\sigma$}\\ q^{\ell(\sigma)+1}(1+q)^{r-h(\mu)-2\ell(\sigma)-1}&\mbox{otherwise.}\end{cases}$ (10)
@@
$\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)=\begin{cases}q^{2+\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)-2}&\mbox{if $s_{i-1}$ and $s_{j+1}$ are not containted in $\sigma$}\\ q^{1+\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)-1}&\mbox{if $s_{i-1}$ is contained in $\sigma$ but $s_{j+1}$ is not}\\ q^{1+\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)-1}&\mbox{if $s_{i-1}$ is not contained in $\sigma$ but $s_{j+1}$ is}\\ q^{\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)}&\mbox{if $s_{i-1}$ and $s{j+1}$are contained in $\sigma$.}\end{cases}$
4 The $q$-Multiplicity of a Positive Root is a Power of $q$
@@
Fix $1\leq i\leq j\leq r$ and let $\mu=\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j}\in\Phi^{+}$. Then
@@
    where $\beta$ is the highest root of $\mathfrak{sl}_{i}(\mathbb{C})$, $\gamma$ is the sum of the fundamental weights of $\mathfrak{sl}_{i}(\mathbb{C})$, and $\sigma_{1}\in\mathcal{A}(\beta,0)$ [6, Theorem 2.1].
@@
    where $\beta^{\prime}$ is the highest root of $\mathfrak{sl}_{r-j+1}(\mathbb{C})$, $\gamma^{\prime}$ is the sum of the fundamental weights of $\mathfrak{sl}_{r-j+1}(\mathbb{C})$, and $\sigma_{2}\in\mathcal{A}(\beta^{\prime},0)$ [6, Theorem 2.1].
    By [6, Proposition 3.2], note that
@@
    where the equality in (25) follows from [2, Theorem 6.1].
@@
    where $\beta^{\prime}$ is the highest root of $\mathfrak{sl}_{r-j}(\mathbb{C})$, $\gamma^{\prime}$ is the sum of the fundamental weights of $\mathfrak{sl}_{r-j}(\mathbb{C})$, and $\sigma_{2}\in\mathcal{A}(\beta^{\prime},0)$ [6, Theorem 2.1].
    By [6, Proposition 3.2], note that
@@
    where expression (27) follows from [2, Theorem 6.1].
@@
    where $\beta$ is the highest root of $\mathfrak{sl}_{i-1}(\mathbb{C})$, $\gamma$ is the sum of the fundamental weights of $\mathfrak{sl}_{i-1}(\mathbb{C})$, and $\sigma_{1}\in\mathcal{A}(\beta,0)$ [6, Theorem 2.1].
@@
    By [6, Proposition 3.2] note that
@@
    where expression (29) follows from [2, Theorem 6.1].
@@
    By [6, Proposition 3.2] note that
@@
    where expression (31) follows from [2, Theorem 6.1].
@@
First, consider $\sum_{\sigma\in{N_{j+1}}}(-1)^{\ell(\sigma)}\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$. By Lemma 3.2, we know the maximum number of $\sigma\in\mathcal{A}(\tilde{\alpha},\mu)$. By Lemma 3.1, we know there are $\binom{r-j-1-k}{k}$ elements $\sigma\in N_{j+1}$ of length $k$ and that the maximum length of any $\sigma\in N_{j+1}$ is $\left\lfloor\frac{r-j-1}{2}\right\rfloor$. Proposition 3.6 gives the value of the $q$-analog $\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$ for $\sigma\in N_{j+1}$. Thus, if $\sigma\in N_{j+1}$ and we use [6, Proposition 3.2], then the first sum in equation (32) becomes
@@
Similarly, consider $\sum_{\sigma\in{Y_{j+1}}}(-1)^{\ell(\sigma)}\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$. By Lemma 3.2, we know there are $\binom{r-j-2-k}{k}$ elements $\sigma\in Y_{j+1}$ of length $k$ and that the maximum length of any $\sigma\in Y_{j+1}$ is $\left\lfloor\frac{r-j-2}{2}\right\rfloor$. Proposition 3.6 gives the value of the $q$-analog $\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$ for $\sigma\in Y_{j+1}$. Thus, if $\sigma\in Y_{j+1}$ and we use [6, Proposition 3.2], then the second sum in equation (32) becomes
@@
First, consider $\sum_{\sigma\in N_{i-1}}(-1)^{\ell(\sigma)}\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$. By Lemma 3.2, we know the maximum length of an element $\sigma\in N_{i-1}$ is $\left\lfloor\frac{i-2}{2}\right\rfloor$. By Lemma 3.1, we know there are $\binom{i-2-k}{k}$ elements $\sigma\in N_{i-1}$. Proposition 3.6 gives the value of the $q$-analog $\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$ for $\sigma\in N_{i-1}$. Thus if $\sigma\in N_{i-1}$ and we use [6, Proposition 3.2], the first sum in equation (36) becomes
@@
Similarly, consider $\sum_{\sigma\in Y_{i-1}}(-1)^{\ell(\sigma)}\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$. By Lemma 3.2, we know the maximum length of an element $\sigma\in Y_{i-1}$ is $\left\lfloor\frac{i-3}{2}\right\rfloor$. By Lemma 3.1, we know there are $\binom{i-3-k}{k}$ elements $\sigma\in Y_{i-1}$. Proposition 3.6 gives the value of the $q$-analog $\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$ for $\sigma\in Y_{i-1}$. Thus if $\sigma\in Y_{i-1}$ and we use [6, Proposition 3.2], the second sum in equation (36) becomes

### v2

@@ 2 Weyl Alternation Sets
We proceed by induction. For $n=1$, the only options are to select no integers or to select $1$. There are two such choices and note $F_{1+2}=F_{3}=2$, and more over $F_{3}=F_{2}+F_{1}$, as claimed. Assume that for all $k\leq n$ the number of ways to select a subset of $[k]$ consisting of only nonconsecutive integers is given by $F_{k+2}$. Now let $k=n+1$. Note that for any subset of $[n-1]$ consisting of nonconsecutive integers, we can include the integer $n+1$ to get a subset of $[n+1]$ consisting of nonconsecutive integers. Then by induction, there are $F_{n-1+2}=F_{n+1}$ many such subsets, all of which contain the integer $n+1$. Moreover, any subset of $[n]$ consisting of nonconsecutive integers is also a subset of $[n+1]$ with the same property and does not contain $n+1$. By induction hypothesis, there are $F_{n+2}$ such subsets. As these subsets are disjoint (since they either contain $n+1$ or not), then we have shown that the number of subsets of $[n+1]$ consisting of nonconsecutive integers is given by $F_{n+2}+F_{n+1}$, which is equal to $F_{n+3}$, as desired. ∎
In the following section, we use the enumerative result Harris gives in [10] which counts subsets of $k$ nonconsecutive integers from the set $[n]$.
@@
Fix $1\leq i\leq j\leq r$ and let $\tilde{\alpha}$ be the highest root of $\mathfrak{sl}_{r+1}(\mathbb{C})$. Then
@@
Selecting a subset consisting of nonconsecutive integers from the set $\{2,\ldots,i-1\}$ is the same as selecting a subset consisting of nonconsecutive integers from the set $\{1,\ldots,i-2\}$. By Proposition 2.3, there are $F_{i}$ such subsets. The same process can be done for selecting a subset consisting of nonconsecutive integers from the set $\{j+1,\ldots,r-1\}$ to get $F_{r-j+1}$ such subsets.
The choices for subsets $\{2,\ldots,i-1\}$ and $\{j+1,\ldots,r-1\}$ are disjoint and independent of each other. Thus, the total number of ways to construct an element of $\mathcal{A}(\tilde{\alpha},\alpha_{i}+\cdots+\alpha_{j})$ is given by the product $F_{i}\cdot F_{r-j+1}$. Therefore, $|\mathcal{A}(\tilde{\alpha},\alpha_{i}+\cdots+\alpha_{j})|=F_{i}\cdot F_{r-j+1}$. ∎
@@ 3 The $q$-Analog of Kostant’s Partition Function
|$|\{\sigma\in\mathcal{A}(\lambda,\mu)\;|\;\ell(\sigma)=k\mbox{ and$\sigma$ contains $s_{j+1}$}\}|$ $=\binom{r-j-2-k}{k}\text{and}$ (6)
@@
|$|\{\sigma\in\mathcal{A}(\lambda,\mu)\;|\;\ell(\sigma)=k\mbox{ and$\sigma$ contains $s_{i-1}$}\}|$ $=\binom{i-3-k}{k}\text{and}$ (8)
@@
Let $\beta=\alpha_{1}+\alpha_{2}+\cdots+\alpha_{s}$ and $\gamma=\alpha_{t}+\alpha_{t+1}+\cdots+\alpha_{s+t-1}$. Notice that $h(\beta)=s$ and $h(\gamma)=(s+t-1)-t+1=s$. By [9, Proposition 3.1], we know
@@
The following propositions give the value of the $q$-analog of Kostant’s partition function when the input is $\sigma(\tilde{\alpha}+\rho)-\rho-\mu$, with $\mu\in\Phi^{+}$ and $\sigma\in\mathcal{A}(\tilde{\alpha},\mu)$. Similar to Proposition 6.1 in [4], the following propositions are closely related. However, they are broken up into different cases based on how $\mu$ is constructed. Thus, we adapt the techniques from [4] in each of the proofs when needed.
@@
    $\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$ =$=\begin{cases}q^{\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)}&\mbox{if $s_{j+1}$ is contained in $\sigma$}\\ q^{1+\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)-1}&\mbox{otherwise.}\end{cases}$ (10)
@@
    $\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$ =$=\begin{cases}q^{\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)}&\mbox{if $s_{i-1}$ is containted in $\sigma$}\\ q^{1+\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)-1}&\mbox{otherwise.}\end{cases}$ (11)
@@
$\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)=\begin{cases}q^{2+\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)-2}&\mbox{if $s_{i-1}$ and $s_{j+1}$ are not containted in $\sigma$}\\ q^{1+\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)-1}&\mbox{if $s_{i-1}$ is contained in $\sigma$ but $s_{j+1}$ is not}\\ q^{1+\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)-1}&\mbox{if $s_{i-1}$ is not contained in $\sigma$ but $s_{j+1}$ is}\\ q^{\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)}&\mbox{if $s_{i-1}$ and $s_{j+1}$ are contained in $\sigma$.}\end{cases}$
4 The $q$-Analog of Kostant’s Weight Multiplicity Formula of a Positive Root is a Power of $q$
@@
Fix $1\leq i\leq j\leq r$ and let $\tilde{\alpha}$ be the highest root of $\mathfrak{sl}_{r+1}(\mathbb{C})$. Then
@@
    where $\beta$ is the highest root of $\mathfrak{sl}_{i}(\mathbb{C})$, $\gamma$ is the sum of the fundamental weights of $\mathfrak{sl}_{i}(\mathbb{C})$, and $\sigma_{1}\in\mathcal{A}(\beta,0)$ [9, Theorem 2.1].
@@
    where $\beta^{\prime}$ is the highest root of $\mathfrak{sl}_{r-j+1}(\mathbb{C})$, $\gamma^{\prime}$ is the sum of the fundamental weights of $\mathfrak{sl}_{r-j+1}(\mathbb{C})$, and $\sigma_{2}\in\mathcal{A}(\beta^{\prime},0)$ [9, Theorem 2.1].
    By [9, Proposition 3.2], note that
@@
    where the equality in (26) follows from [4, Theorem 6.1].
@@
    where $\beta^{\prime}$ is the highest root of $\mathfrak{sl}_{r-j}(\mathbb{C})$, $\gamma^{\prime}$ is the sum of the fundamental weights of $\mathfrak{sl}_{r-j}(\mathbb{C})$, and $\sigma_{2}\in\mathcal{A}(\beta^{\prime},0)$ [9, Theorem 2.1].
    By [9, Proposition 3.2], note that
@@
    where expression (28) follows from [4, Theorem 6.1].
@@
    where $\beta$ is the highest root of $\mathfrak{sl}_{i-1}(\mathbb{C})$, $\gamma$ is the sum of the fundamental weights of $\mathfrak{sl}_{i-1}(\mathbb{C})$, and $\sigma_{1}\in\mathcal{A}(\beta,0)$ [9, Theorem 2.1].
@@
    By [9, Proposition 3.2] note that
@@
    where expression (30) follows from [4, Theorem 6.1].
@@
    By [9, Proposition 3.2] note that
@@
    where expression (32) follows from [4, Theorem 6.1].
@@
First, consider $\sum_{\sigma\in{N_{j+1}}}(-1)^{\ell(\sigma)}\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$. By Lemma 3.2, we know the maximum number of $\sigma\in\mathcal{A}(\tilde{\alpha},\mu)$. By Lemma 3.1, we know there are $\binom{r-j-1-k}{k}$ elements $\sigma\in N_{j+1}$ of length $k$ and that the maximum length of any $\sigma\in N_{j+1}$ is $\left\lfloor\frac{r-j-1}{2}\right\rfloor$. Proposition 3.6 gives the value of the $q$-analog $\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$ for $\sigma\in N_{j+1}$. Thus, if $\sigma\in N_{j+1}$ and we use [9, Proposition 3.2], then the first sum in equation (33) becomes
@@
Similarly, consider $\sum_{\sigma\in{Y_{j+1}}}(-1)^{\ell(\sigma)}\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$. By Lemma 3.2, we know there are $\binom{r-j-2-k}{k}$ elements $\sigma\in Y_{j+1}$ of length $k$ and that the maximum length of any $\sigma\in Y_{j+1}$ is $\left\lfloor\frac{r-j-2}{2}\right\rfloor$. Proposition 3.6 gives the value of the $q$-analog $\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$ for $\sigma\in Y_{j+1}$. Thus, if $\sigma\in Y_{j+1}$ and we use [9, Proposition 3.2], then the second sum in equation (33) becomes
@@
First, consider $\sum_{\sigma\in N_{i-1}}(-1)^{\ell(\sigma)}\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$. By Lemma 3.2, we know the maximum length of an element $\sigma\in N_{i-1}$ is $\left\lfloor\frac{i-2}{2}\right\rfloor$. By Lemma 3.1, we know there are $\binom{i-2-k}{k}$ elements $\sigma\in N_{i-1}$. Proposition 3.6 gives the value of the $q$-analog $\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$ for $\sigma\in N_{i-1}$. Thus if $\sigma\in N_{i-1}$ and we use [9, Proposition 3.2], the first sum in equation (37) becomes
@@
Similarly, consider $\sum_{\sigma\in Y_{i-1}}(-1)^{\ell(\sigma)}\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$. By Lemma 3.2, we know the maximum length of an element $\sigma\in Y_{i-1}$ is $\left\lfloor\frac{i-3}{2}\right\rfloor$. By Lemma 3.1, we know there are $\binom{i-3-k}{k}$ elements $\sigma\in Y_{i-1}$. Proposition 3.6 gives the value of the $q$-analog $\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$ for $\sigma\in Y_{i-1}$. Thus if $\sigma\in Y_{i-1}$ and we use [9, Proposition 3.2], the second sum in equation (37) becomes